OpenStudy (kkrpbra):

Write each decimal as a fraction in simplest form 1.) -2.05 2.) 3.502 3.) 0.064 4.) -8.4 5.) 0.004 6.) -0.70 7.) 5.25 8.) 0.128 9.) 0.0004

OpenStudy (anonymous):

I'll do the first few. For (1) and (2), I'll expand each decimal place as its own fraction. 1) -2.05 = -2 - 1/20 = -40/20 - 1/20 = -41/20 2) 3.502 = 3 + 0.5 + 0.002 = 3 + 1/2 + 1/500 = 1500/500 + 250/500 + 1/500 = 1751/500 For (3), I'll write it directly as a fraction. 3) 0.064 = 64/1000 = 32/500 = 16/250 = 8/125 Same way with (4) 4) -8.4 = -84/10 = - 42*2/(5*2) = -42/5